OpenStudy (anonymous):

Find the slope of the tangent line to the graph of the function f(x)=x e^x at x=7 correct to two decimal places.

sam (.sam.):

dy/dx = xe^x + e^x sub x=7 into equation, dy/dx = 8773

OpenStudy (anonymous):

f'(x) = e^x (x+1) f'(7) = e^7 (7+1) = 8e^7 = 8773.07

OpenStudy (anonymous):

sam, what rule did u use to get the + e^x?